Sol Summary Report – 26 April 2018

Sol 4

Summary Title: Broken Glasses and Martian Pizza

Author: David Attig, Crew Engineer

Mission Status: Super Tremendous Nominal.

Sol Activity Summary: The day started off early with quick individual breakfasts in order to prepare for EVA 8 which left at 0730.

The crew feasted on some homemade pizza, with some fresh salad and soup made with our greenhouse produce. We spent time on personal work and research projects before EVA 9.

While EVA 9 was underway, the HAB was vacuumed and progress made on experiments.

Look Ahead Plan: Tomorrow we will stick close to the Hab due to predicted hot weather and high winds in the afternoon. We will have two EVAs to practice on ATVs, and will spend the afternoon in the Hab practicing medical emergency situations and reviewing fire safety.

Anomalies in work:

1. There was a pair of sunglasses lost during the afternoon EVA. They were found, but had been driven over by a Rover.

2. There was a missed picture opportunity; the crew member who was asked to take it did a poor job and the crew member who asked for it did not review the picture before returning from EVA.

Weather: Clear skies in AM, chilly in morning and quickly warming during the day. Warm and breezy in PM.

Crew Physical Status: Nominal

EVA:

EVA 8 – went well, gray moon area was explored. There is incredible scenery, the landscape matches up with its name very well. Upon returning, they brought materials in from the RAM to work on in the HAB.

EVA 9 – Saw a fantastic view of Candor Chasma.

Sol Summary Report April 25th

Crew 193 Sol Summary Report – 25 APR 2018

Sol 3

Summary Title: Life Will Find a Way

Author: Doug Campbell, Executive Officer and Science Officer

Mission Status: Sub-optimal due to minor crew health issues but expected to be back to nominal tomorrow

Sol Activity Summary: The day started off early with quick individual breakfasts in order to prepare for EVA 6 which left at 0730. EVA 6 was a successful failure in that valuable exploration of Lith Canyon was completed; however, the intended destination was the White Moon area! While EVA 6 was occurring, the plants were watered in the Green Hab and some vegetables were harvested. Our crew astronomer made a quick trip to the observatory and did some troubleshooting with the GPS units. Our crew engineer did work around the hab and communicated with mission control.

The crew lunched together (thanks to our commander for cooking!) and then spent time on personal work and research projects before EVA 7. EVA seven was cut short at 1600 by the Health and Safety Officer over concerns of crew dehydration. Late afternoon was spent on reports. The plan for tonight is to watch a movie as a team, perhaps…. The Martian!!

Look Ahead Plan: EVA 8 will attempt to complete the mission plan that was submitted for EVA 6 but not completed. EVA 9 will be exploring further into Candor Chasma as part of the GPS mapping study. We will continue to work on learning ASL for 5 minutes each day as well as sharing stories about our lives so that the crew gets to know each other better.

Anomalies in work:

1. Radio #6 has a poor battery which loses charge quickly. It caused Gold 001 to be without communication during EVA 6. This radio has been put out of service by crew 193
2. EVA 7 was cut short due to health concerns over crew dehydration

Weather: Clear skies in AM, partly cloudy and gusty in PM.

Crew Physical Status: Some blisters, one incidence of crew mild dehydration. All members are resting and back to nominal by days end.

EVA:
EVA 6 – Successfully explored the edges and the base of Lith Canyon. As per the title of this report, the EVA team found group of fully blossomed wild flowers on the edge of a large, sandy rock formation with no other flora or fauna in sight; proving that even on mars, life will find a way.
EVA 7 – Began the EVA in the RAM to paint the new sign for the MDRS campus. The crew then proceeded to climb the hill up to hab ridge. However, during the climb, one of the members began to not feel well and the remainder of the EVA was scrubbed.

Reports filed:
GreenHab
Science Report
Astronomy Report
Journalist’s Report
Photo & Photo of the Day Report

Reports to be filed:
Operations Report
EVA Report

Support Requested:

Additional fuel for ATVs.
Orange or green paint for the MDRS sign (or both)
Hot pads so we can put hot cookware on the table without wrecking the tabletop.

Sol 2 Summary – April 24th

Crew 193 Sol Summary Report 24APR2018

Sol 2

Summary Title: Every Sol We Hustlin’…

Author: Dr. Shawna Pandya, Crew Health and Safety Officer

Mission Status: Nominal

Sol Activity Summary: EVAs 4 & 5 complete, focusing on science, exploration, and sign-work. In-hab activities consisted of usual maintenance, operations and science. As a team, we managed to fit in some bonding and learn some basic ASL (American Sign Language) phrases from our deaf crewmate. Crucially, I learned to say, “I NEED MY COFFEE!!” in American Sign Language.

Look Ahead Plan: EVA 6 will plan for refueling the ATVs, then North-ward exploration along ‘The Moon’ – objectives: exploration; visor fogging study; emergency shelter scouting on EVA study; crew comfort, health, well-being and safety on EVA study; GPS accuracy on EVA. EVA 7 will plan on continued work in the RAM on the new MDRS sign, as well as visor fogging and crew comfort studies. In-hab activities will focus on the usual maintenance, operations and science. Time permitting, we are for some more crew training in ASL.

Anomalies in work:

1)Team on EVA 5 was forced to use 1 Rover, as the Beartooth communication with our deaf crewmate failed, and the LCD board was taken instead, and could not be safely transported in the ATV.

2)Will be decommissioning spacesuit 7, as the visor is deemed too scratched to allow for safe visualization.

3)One of the two-piece spacesuits came apart at the mechanical neck-ring/backpack interface while on EVA. It will be decommissioned.

Weather: Clear skies in AM, partly cloudy and gusty in PM.

Crew Physical Status: Some blisters, otherwise nominal.

EVA:

EVA 4 – Southeast exploration along Phobos Peak and Kissing Camels’ Ridge E & W; continued work on scouting, crew comfort, visor fogging studies.

EVA 5 – Measurements taken for second MDRS sign; East-ward exploration along Tank Road; continued work on scouting, crew comfort, visor fogging studies.

Reports filed:

GreenHab
EVA 4 Report
Science Report
Astronomy Report

Reports to be filed:
Commander or Journalist’s Report
Operations Report
EVA 5 Report
Photo & Photo of the Day Report

Support Requested:

-Additional fuel for ATVs.

-Requesting location of dropship of supplies that include Vicks Lozenges.

-Happy to attempt to replace the visor on spacesuit 7 if Mission Support can provide us with the appropriate procedure and supplies.

-Happy to attempt to repair the two-piece spacesuit if Mission Support can provide us with the appropriate procedure and supplies.

Sol Summary – April 23rd

Crew no. 193 Sol Summary Report 23/04/2018

Sol: 1

Learning the Ropes

Eric Shear

Mission Status: Everything nominal

Sol Activity Summary: Carried out two EVAs, including one extended EVA and an emergency 1-person EVA. Monitored the Sun for sunspot activity

Look-Ahead Plan: We continue our exploration of Mars. Our EVA 4 will be planned around a new project “GPS Route Measurements”

Anomalies in work: We had a stranded astronomer whom we rescued. We fixed a signpost with the tools available at the end of our EVA in an extended EVA. Also, there was an emergency 1-person EVA to secure the RAM Module, with no issues.

Weather: Very clear and hot in morning/midday, cooler and cloudy in afternoon

Crew Physical Status: All in good health

EVA 1:

Went to Miracle Ritual, scouted locations for emergency evacuation, conducted an EVA health/wellness study, and a visor fogging study.

EVA 2:

Went to Miracle Ritual, scouted locations for emergency evacuation, conducted the rest of the health/wellness study, executed an emergency assist of the telescope repairman, and repaired the road sign to the hab.

Emergency EVA:

1 person exited the hab to secure the Repair and Assembly Module, which door had been left open on the previous EVA. There were no issues.

Reports to be filed:

Commander’s Report
Operations Report
Greenhouse Report
EVA Reports
Astronomy Report

Support Requested: We need aluminum foil and dish scrubbing pads (the hard ones for tough stains and grease)

Sol Summary – April 21st

Crew 192 Sol Summary Report 21Apr2018

Sol 13

Summary Title: Prepare for Departure

Author: Victoria Varone

Mission Status: Nominal

Sol Activity Summary: We welcomed Crew 193 to the hab and gave them an introduction and orientation to Mars life.

Look Ahead Plan: Departure and adjusting to life back on Earth.

Anomalies in Work: None

Weather: Beautiful

Crew Physical Status: All crew are physically healthy and safe.

EVA: None

Reports to be filed: Sol Summary Report, Operations Report, GreenHab Report

Support Requested: Please remove Joe, Ashok, and Andreea from Mission Support emails. Keep Victoria and Richard on for the duration of Crew 193’s rotation.

Maintenance request: New lock for Musk Observatory and fluorescent lights for lower deck of hab.

Sol Summary – April 20th

Crew 192 Sol Summary Report 20Apr2018

Sol 12

Summary Title: Splash Down

Author: Victoria Varone

Mission Status: Nominal

Sol Activity Summary: The crew divided up our last day in sim into two fantastic EVAs and a mission to clean the hab from top to bottom for the incoming crew and close out all of our reports. We went through each of the buildings in the hab site and made sure everything is clean, organized, and in great shape for use by future crews.

Look Ahead Plan: Finishing up our last reports and inventories to prepare for incoming crew and our departure.

Anomalies in Work: None

Weather: Cloudy and rainy in the morning, warmer and sunny in the afternoon.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, Photo Report, Commander’s Report, GreenHab Report, Astronomy Report, End of Mission Report, Science Report

Support Requested: Note: Filter for loft water tank is already light brown/orange in color.

Maintenance request: Lock for Musk Observatory is no good anymore, it was severely jammed and we tried various fixes but it still sticks. Recommend replacement. Also we will need fluorescent light bulbs soon, They are flickering. Finally had to secure the water under the sink in the upstairs sink because the leak has gotten worse.

Sol Summary – April 19th

Crew 192 Sol Summary Report 19Apr2018

Sol 11

Summary Title:

Author: Victoria Varone

Mission Status: Nearing it’s completion, we’ve begun field day work to present the hab in the best possible condition for the next crew.

Sol Activity Summary: A wonderful morning EVA across Hab Ridge Road capped off our intense EVAs for our mission, where we hiked along the beautiful ridge and collected some data on helmet fogging inside our spacesuits.

We climbed down and took some great photos of the beautiful sites and structures on our way back to the hab.

Upon returning, we had lunch and then executed a short engineering EVA to finish up some close-out tasks for the end of our mission and the incoming Crew 193.

Look Ahead Plan: Planning a few wrap up EVAs to collect last bit of Data. Getting the hab ready for Crew 193.

Anomalies in Work: None

Weather: Cool in the morning, then warm and sunny, ending with clouds in the afternoon.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, Photo Report, Commander’s Report, GreenHab Report, Journalist Report

Support Requested: None

Maintenance request: See Operations Report

Sol Summary – April 18th

Crew 192 Sol Summary Report 18Apr2018

Sol 10

Summary Title: Canyons, Baby

Author: Victoria Varone

Mission Status: A beautiful day for the crew, mission nearing its end.

Sol Activity Summary: The crew went on two amazing EVAs today, exploring some of the major Martian canyons in the area. Lith Canyon and Murphy Canyon provided some of the most breathtaking views we’ve seen so far.

Then we get a little sun-gazing in at the Musk Observatory before cooking an awesome dinner for the crew that was a wonderful example of a team effort.

Look Ahead Plan: Aiming to make our last few EVAs fantastic ones and get in some great stargazing before we leave.

Anomalies in Work: None

Weather: Freezing in the morning, a bit warmer throughout midday and the afternoon. Calm with little wind and partly cloudy.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report, Astronomy Report

Support Requested: Please add Doug Campbell, incoming with Crew 193, to the Mission Support list

Maintenance request: See Operations Report

Sol Summary – April 17th

Crew 192 Sol Summary Report 17Apr2018

Sol 9

Summary Title: Don’t Feed the Penguin

Author: Victoria Varone

Mission Status: Nearing its end, aiming to fit in as much exploration as possible.

Sol Activity Summary: The crew went on two great EVAs today, exploring parts of the Martian terrain that we haven’t yet seen. It’s fascinating to see just how many different kinds of soil, rock, and sediment you can find even in just one area, in varying colors, textures, shapes, and sizes. The winds were strong today (and last night) so extra caution was exercised during EVAs.

Look Ahead Plan: Trying to fit in as many EVAs as possible (pending weather) since weekend weather may not be suitable and our mission ends on Saturday.

Anomalies in Work: None

Weather: Chilly, mix of clouds and sun, strong winds throughout the night and most of the day.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report

Support Requested: None

Maintenance request: See Operations Report

Sol Summary – April 16th

Crew 192 Sol Summary Report 16Apr2018

Sol 8

Summary Title: Houston, We Have a [Mouse] Problem

Author: Victoria Varone

Mission Status:

Sol Activity Summary: Our day started with a visit from a tiny companion ‒ code name “Paul Watney” in honor of our missing crew member and “The Martian” ‒ who made his way into the upper deck of the hab. He accompanied us in the hab for the day but then decided to return to Mars. (see Photo Report)

Our first EVA of the day was a long walking hike along the ridge. We climbed some of the dunes and hills and had to get creative along the way as there weren’t outlets or shallow spots where we thought they would be. One of the highlights of the hike was making our way up to the hab overview, which offered some amazing Martian views overlooking the hab.

Our second EVA was shortened due to strong winds that increased as the day went one. It was turned into a short relocation EVA where we dropped off our tiny friend “Paul Watney” back onto the vast plains of Mars.

The evening consisted of some more cooking for the crew, baking peanut butter cookies, and watching “Apollo 13.”

Look Ahead Plan: Finish cleaning Water Tank and final skirt HAB clean up. Exploration continues to the south and spreading North.

Anomalies in Work: None

Weather: Chilly early on, warmer later in the day, mostly cloudy all day.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report

Support Requested: None

Maintenance request: See Operations Report